OpenStudy (anonymous):

Subtract the following fractions 4 3/5 - 1 2/3 (simplify if needed)

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@Christos

OpenStudy (anonymous):

@Maddog1456

OpenStudy (anonymous):

ok

OpenStudy (maddog1456):

fractions aint my strong suit

OpenStudy (maddog1456):

sorry

OpenStudy (anonymous):

okay

OpenStudy (flower2468):

wait sorry you need to make them into improper fractions SO sorry

OpenStudy (flower2468):

so improper fractions 4*5+3=23 on top and 1*3+1=4 on top

OpenStudy (anonymous):

it's okay

OpenStudy (anonymous):

okay thank you! I will give you a medal!

OpenStudy (flower2468):

Then you need to find common denominators

OpenStudy (anonymous):

ok

OpenStudy (flower2468):

5*3=15 and 23*3=69 or 69/15 and then 3*5=15 and 4*5=20 or 20/15

OpenStudy (flower2468):

then subtract 69-20 and you get....

OpenStudy (anonymous):

49

OpenStudy (flower2468):

okay then the fraction is 49/15 and then we need to simplify it if needed

OpenStudy (anonymous):

yes

OpenStudy (flower2468):

okay then how many times can 49 go into 15

OpenStudy (flower2468):

PINK?